OpenStudy (clamin):

What is the 11th term of the following geometric sequence? 7/9, -7/3,  7, -21, 63

OpenStudy (clamin):

The choices are  a. 15,309 b. -15, 309 c. -45, 927 d. 45,927

Nnesha (nnesha):

do you know how to find common ration or what is common ratio ?

OpenStudy (clamin):

The gap of two numbers

Nnesha (nnesha):

to find common ratio you can divide next term with the previous one \[\large\rm r=\frac{ a_2 }{ a_1} ~~ or~~\frac{ a_3 }{ a_2 } ~~or ~~\frac{ a_4 }{ a_3 }\] then use the general form of geometeric sequence \[\huge\rm a_n = a_1 \times (r)^{n-1} \] a_1 = first term n = number of term that u have to find r= ratio